Can someone please help me decipher this question.

The return from investing in foreign assets is affected by the performance of the foreign assets and the fluctuation in the exchange rate of the foreign country.  Using the information give below, calculate the rate of dollar return from investing in British Treasury bond with face value of £100 if you bought it year ago and sold it today:

                        Price of British bond year ago:              £101

                        Price of British bond today:                   £103

                        British pound exchange rate last year:   $1.2800/£

                        British pound exchange rate today:        $1.2950/£
Answer #1

Last year while conversion to buy a bond with price of £101 and exchange rate of $1.2800/£, I have to invest =1.28*101=$129.28

Now the price of bond increased to £103, so now if we exchange back £103 in to $ at exchange rate $1.2950/£ = 103*1.295 = $133.385

rate of dollar return in 1 year = (amount received - initial investment)/initial investment = (133.385 - 129.28)/129.28= 3.2%